### v2

@@ 1 Introduction
In model-driven software and systems engineering, models serve as the primary artifacts for development, evolving throughout their lifecycle due to modifications, refinements, and refactorings. To facilitate a less error-prone development process and ensure the preservation of critical model properties, automated model analyses, such as semantic differencing, can be employed [1].
Semantic differencing is a comparative model analysis technique that evaluates two models written in the same language by examining their legal instances as defined by a language-specific formal semantics [2]. A semantic difference is identified when a "?diff-witness? exists—an instance that is valid in one model but not in the other. Conversely, if no such witness is found, the first model can be considered a semantic refinement of the second. Various semantic difference operators have been developed for different modeling constructs, including activity diagrams [3, 4], class diagrams [5, 6], feature models [7], variants of statecharts [8, 9, 10], and sequence diagrams [11].
Despite the extensive exploration of semantic differencing in static structural models, such as class diagrams, and isolated behavioral models, like statecharts, component-and-connector architectures present unique challenges due to their dynamic nature and compositional complexity. Unlike static object structures that do not define dynamic behavior, component-and-connector models necessitate the consideration of both individual component behaviors and their interactions within compositions. This paper addresses these challenges by focusing on semantic differencing specifically for MontiArc models [12, 13, 14], a topic that has yet to be comprehensively covered in existing literature.
@@
To analyze behavioral differences in component and connector architectures, we aim to leverage Dynamic Symbolic Execution (DSE) [15], a variant of symbolic execution [16], to develop a semantic differencing operator specifically for MontiArc models. Building upon the fundamental semantic-difference framework established in previous publications, we apply DSE as a novel technique for identifying semantic differences in these models.
MontiArc [12, 13, 14] is an architecture modeling framework developed with the MontiCore [^1] https://monticore.github.io/monticore/ language workbench [17, 18] that offers an architecture description language for designing both software and cyber-physical systems [19]. It enables modular hierarchical modeling [20] of architectural components and their message-oriented, asynchronous communication. The input-output behavior of atomic components can then be specified by a variant of Statecharts (SCs) [21, 22] based on the FOCUS formalism [23].
To facilitate simulations, MontiArc includes a code generator that produces executable Java code [13, 12]. We have extended this generator to create symbolic code suitable for DSE. This symbolic code can be invoked using a newly developed tool designed to initiate the analysis. A default execution strategy for DSE is provided, with support for additional user-defined strategies. These strategies utilize execution data from the MontiArc model and employ Z3 [24] as a satisfiability modulo theories (SMT) solver.
@@ 2 Dynamic Symbolic Execution
Dynamic Symbolic Execution (DSE) is a variant of symbolic execution that generates concrete inputs alongside symbolic inputs [15]. In symbolic execution, each input is represented as a symbolic constant. Each execution path of a program is associated with a boolean expression known as the path condition. These path conditions are initialized with the value true and define the properties that must be satisfied by a concrete value in order to execute that specific path. During the process of symbolic execution, an execution tree is constructed, representing all paths identified by the chosen execution strategy [16]. Concrete values can be derived as inputs by solving individual path constraints using a Satisfiability Modulo Theories (SMT) solver.
@@
Symbolic execution has inherent limitations, particularly in cases where certain functions cannot be effectively executed symbolically, such as cryptographic functions [25]. Dynamic Symbolic Execution (DSE), also known as concolic execution, addresses these limitations by combining both concrete and symbolic execution of a program, executing it with both concrete and symbolic values simultaneously [26, 27, 28, 29].
@@ 3 Related Work
Symbolic execution is a powerful technique used in program analysis that systematically explores program paths by treating inputs as symbolic variables. However, it is computationally intensive, and while advancements in modern computer hardware have facilitated the development of efficient theorem provers and scalable analysis tools, the challenges inherent in symbolic execution have prompted renewed interest in this area [25].
Despite its strengths, symbolic execution faces two primary limitations: path explosion and the complexity of constraint solving. These challenges are also prevalent in Dynamic Symbolic Execution (DSE) and significantly hinder scalability. For instance, constructs such as loops can lead to an exponential increase in the number of feasible execution paths, making it infeasible to systematically analyze all possible paths in large systems or programs. Additionally, the second limitation pertains to constraint solving, which is an NP-hard problem. In larger systems, path constraints can involve intricate combinations of conditions, further complicating the analysis. Consequently, achieving complete path coverage remains elusive in practice [28, 30, 29, 31].
@@
CUTE [27], a concurrent unit testing engine, can efficiently explore paths in C code, achieving high branch coverage and bug detection. Due to the limitations of DSE with respect to path explosion, an execution strategy has been used to counteract this. CUTE uses a bounded depth-first search to counteract an infinite exhaustive search of the entire computation tree. Three optimizations have been implemented to counteract the limitations of constraint solving: first, the fast unsatisfiability check, which checks whether the last constraint of the path condition is syntactically the negation of any previous constraints, second, common sub-constraints elimination, in which the solver identifies and eliminates common arithmetic sub-constraints, and third, incremental solving, where the solver identifies dependencies between sub-constraints and exploits them to solve constraints faster and keep solutions similar. These optimizations reduce the number of sub-constraints and thus optimize the runtime.
@@
jCUTE [32] was one of the pioneering symbolic execution tools designed specifically for Java programs. However, as jCUTE is no longer actively maintained, JDart has emerged as a robust replacement [33]. JDart employs DSE to analyze Java programs, primarily focusing on assertion checking. Through this approach, JDart can either identify assertion violations, exhaustively explore all program paths, or reach resource limits during analysis.
@@
One of the key advantages of JDart is its capability to handle complex software systems, as exemplified by its development to analyze intricate NASA software. JDart’s architecture is modular, comprising two primary components: the executer and the explorer. The executer is responsible for executing the program and collecting symbolic constraints, leveraging the Java PathFinder framework [34, 35]. In contrast, the explorer determines the search and execution strategy employed during analysis.
@@
JDart was designed with two main objectives in mind. The first is to create a robust framework capable of managing industrial software challenges, such as handling crashes and addressing the constraints of constraint solving. The second objective is to establish a modular and extensible platform, allowing for the interchangeability of components, the use of various constraint solvers, and the implementation of multiple search strategies or termination conditions [36].
@@
SAGE [30], developed by Microsoft, is an automated white-box fuzz testing tool widely utilized for detecting bugs and vulnerabilities in applications, such as those running on Windows. SAGE implements a technique known as Generational Search, which minimizes redundancy while maximizing the generation of new test cases. This systematic approach allows for effective analysis of the state space, enabling SAGE to manage large applications with substantial input sizes. Additionally, SAGE employs heuristics to enhance code coverage, further improving its effectiveness in identifying potential issues.
@@
Another tool developed by Godefroid is DART [28], which focuses on automatic software testing using DSE. Building on DART’s capabilities, SMART [25] introduces the concept of compositional analysis, wherein composite functions are decomposed into atomic functions for individual analysis. The results of these analyses are then synthesized through the use of pre- and post-conditions.
@@
A key aspect that differentiates our research from previous work on DSE is our intended goal: to ascertain the semantic differences between executable component-and-connector architecture models. As noted in section˜1, semantic differencing operators have been developed for various modeling languages [37, 5, 7, 10, 11], and component-and-connector structures, such as statechart systems (SCS) or automata, are no exception [8, 9]. However, these approaches typically rely on translations to Büchi automata, which necessitate a finite state space and a defined input-output alphabet. Additionally, to compare entire architectures, syntactic composition of the automata is required, posing challenges when feedback loops are present. Our DSE-based approach, in contrast, circumvents these limitations.
@@ 4 Running Example
MontiArc [12, 13, 14] is an architectural description language specifically designed for component-and-connector architectures. In this framework, components can either be atomic or consist of sub-architectures, allowing for a hierarchical organization of system elements. Communication between these components is achieved through a message-oriented approach, utilizing channels that connect output and input ports. The input-output behavior of atomic components is described by automata that are a variant of SCs [21, 22] based on the FOCUS formalism [23].
@@ 5.1 Major Design Decisions
The second reason for choosing DSE is that not all functions can be effectively analyzed using traditional symbolic execution. For instance, cryptographic functions, such as hash functions, cannot be evaluated through symbolic execution without undermining their security properties. In the case of DSE, these functions can be analyzed by utilizing concrete values for variables. By substituting concrete values, we can evaluate the function and consequently derive new input values for further analysis [25]. In summary, DSE was selected because it enables the analysis of functions that are not amenable to traditional symbolic execution and avoids exploring paths with unsatisfiable conditions.
@@
Z3 is an award-winning SMT solver developed by Microsoft. It is primarily used for predicate abstraction, advanced static checking, and test case generation. Interaction with Z3 can occur through the SMT-LIB format as well as various APIs [24]. Our implementation utilizes the Java API for communication with the solver. Z3 supports multiple theories, including equality, arithmetic operations, and uninterpreted functions. Additionally, it generates a model containing concrete values for all defined constants within the formula. Z3 also includes numerous optimizations and tactics to enhance its performance [38].
@@
Unlike other DSE engines that perform Dynamic Symbolic Execution by executing the program under test and collecting symbolic information without explicit code generation [30, 33, 27], MontiArc employs a code generator that produces executable Java code [13]. In this paper, we evaluate a DSE approach based on this existing code generator and extend it to generate symbolic code. We also discuss alternative methods, such as using an interpreter, in section˜7.
@@ 5.3 Computation of Semantic Differences Between MontiArc Models
Figure 7: Example of semantic difference between semDiffSmallModel and StudentVote, input length 3
When assessing the semantic difference between two models, it is crucial to consider certain specific scenarios. One such scenario arises when a component in the second model contains an automaton that is deemed partial concerning the possible inputs of the first model. In this case, the partial automaton will respond to any unknown input by ignoring it—meaning it neither produces a new output nor transitions to a different state [39]. Consequently, the automaton retains its previous state and output. However, if the output ports are delayed, Null values may be generated. It is essential to address these cases within the automaton’s definition of the component behavior.
@@ 6.2 Results of the Evaluation
Figure 8: Results regarding the smallModel based on input length 4
@@
In the subsequent analysis, we compute the semantic difference using the Path Coverage Controller with a garbage collection trigger. The results of the semantic difference analysis between StudentVote and smallModelSemDiff for varying input lengths are presented in table˜2.
@@
Table 2: Results of the computation of semantic differences between smallModel and semDiffSmallModel
@@ 7 Discussion
Another avenue for enhancing the tool is through decomposition analysis of models, as introduced by  Godefroid [25]. This technique involves breaking down the program under analysis into individual functions. The results from analyzing these functions are then synthesized, taking into account their pre- and post-conditions, to evaluate composite programs. This methodology should be applicable to MontiArc models by examining their atomic components, which are already represented by atomic model artifacts. Implementing this approach will necessitate adjustments to the tool architecture.

### v3

@@ 1 Introduction
In model-driven software and systems engineering, models serve as the primary artifacts for development, evolving throughout their lifecycle due to modifications, refinements, and refactorings. To facilitate a less error-prone development process and ensure the preservation of critical model properties, automated model analyses, such as semantic differencing, can be employed [MRR10].
Semantic differencing is a comparative model analysis technique that evaluates two models written in the same language by examining their legal instances as defined by a language-specific formal semantics [HR04]. A semantic difference is identified when a "?diff-witness? exists—an instance that is valid in one model but not in the other. Conversely, if no such witness is found, the first model can be considered a semantic refinement of the second. Various semantic difference operators have been developed for different modeling constructs, including activity diagrams [MRR11d, KR18], class diagrams [MRR11b, RRS23], feature models [DKMR19], variants of statecharts [BKRW17, BKRW19, DEKR19], and sequence diagrams [Kau21].
Despite the extensive exploration of semantic differencing in static structural models, such as class diagrams, and isolated behavioral models, like statecharts, component-and-connector architectures present unique challenges due to their dynamic nature and compositional complexity. Unlike static object structures that do not define dynamic behavior, component-and-connector models necessitate the consideration of both individual component behaviors and their interactions within compositions. This paper addresses these challenges by focusing on semantic differencing specifically for MontiArc models [HRR12, Hab16, BKRW17a], a topic that has yet to be comprehensively covered in existing literature.
@@
To analyze behavioral differences in component and connector architectures, we aim to leverage Dynamic Symbolic Execution (DSE) [Cadar], a variant of symbolic execution [king], to develop a semantic differencing operator specifically for MontiArc models. Building upon the fundamental semantic-difference framework established in previous publications, we apply DSE as a novel technique for identifying semantic differences in these models.
MontiArc [HRR12, Hab16, BKRW17a] is an architecture modeling framework developed with the MontiCore [^1] https://monticore.github.io/monticore/ language workbench [GKR+06, HKR21] that offers an architecture description language for designing both software and cyber-physical systems [GRSS12]. It enables modular hierarchical modeling [BR07] of architectural components and their message-oriented, asynchronous communication. The input-output behavior of atomic components can then be specified by a variant of Statecharts (SCs) [statechart, Rum16] based on the FOCUS formalism [BS12].
To facilitate simulations, MontiArc includes a code generator that produces executable Java code [Hab16, HRR12]. We have extended this generator to create symbolic code suitable for DSE. This symbolic code can be invoked using a newly developed tool designed to initiate the analysis. A default execution strategy for DSE is provided, with support for additional user-defined strategies. These strategies utilize execution data from the MontiArc model and employ Z3 [efficientZ3] as a satisfiability modulo theories (SMT) solver.
@@ 2 Dynamic Symbolic Execution
Dynamic Symbolic Execution (DSE) is a variant of symbolic execution that generates concrete inputs alongside symbolic inputs [Cadar]. In symbolic execution, each input is represented as a symbolic constant. Each execution path of a program is associated with a boolean expression known as the path condition. These path conditions are initialized with the value true and define the properties that must be satisfied by a concrete value in order to execute that specific path. During the process of symbolic execution, an execution tree is constructed, representing all paths identified by the chosen execution strategy [king]. Concrete values can be derived as inputs by solving individual path constraints using a Satisfiability Modulo Theories (SMT) solver.
@@
Symbolic execution has inherent limitations, particularly in cases where certain functions cannot be effectively executed symbolically, such as cryptographic functions [Godefroid_SMART]. Dynamic Symbolic Execution (DSE), also known as concolic execution, addresses these limitations by combining both concrete and symbolic execution of a program, executing it with both concrete and symbolic values simultaneously [Xiao, Sen, Godefroid_DART, Baldoni].
@@ 3 Related Work
Symbolic execution is a powerful technique used in program analysis that systematically explores program paths by treating inputs as symbolic variables. However, it is computationally intensive, and while advancements in modern computer hardware have facilitated the development of efficient theorem provers and scalable analysis tools, the challenges inherent in symbolic execution have prompted renewed interest in this area [Godefroid_SMART].
Despite its strengths, symbolic execution faces two primary limitations: path explosion and the complexity of constraint solving. These challenges are also prevalent in Dynamic Symbolic Execution (DSE) and significantly hinder scalability. For instance, constructs such as loops can lead to an exponential increase in the number of feasible execution paths, making it infeasible to systematically analyze all possible paths in large systems or programs. Additionally, the second limitation pertains to constraint solving, which is an NP-hard problem. In larger systems, path constraints can involve intricate combinations of conditions, further complicating the analysis. Consequently, achieving complete path coverage remains elusive in practice [Godefroid_DART, Godefroid_Fuzz, Baldoni, Williams].
@@
CUTE [Sen], a concurrent unit testing engine, can efficiently explore paths in C code, achieving high branch coverage and bug detection. Due to the limitations of DSE with respect to path explosion, an execution strategy has been used to counteract this. CUTE uses a bounded depth-first search to counteract an infinite exhaustive search of the entire computation tree. Three optimizations have been implemented to counteract the limitations of constraint solving: first, the fast unsatisfiability check, which checks whether the last constraint of the path condition is syntactically the negation of any previous constraints, second, common sub-constraints elimination, in which the solver identifies and eliminates common arithmetic sub-constraints, and third, incremental solving, where the solver identifies dependencies between sub-constraints and exploits them to solve constraints faster and keep solutions similar. These optimizations reduce the number of sub-constraints and thus optimize the runtime.
@@
jCUTE [jCute] was one of the pioneering symbolic execution tools designed specifically for Java programs. However, as jCUTE is no longer actively maintained, JDart has emerged as a robust replacement [jDart20]. JDart employs DSE to analyze Java programs, primarily focusing on assertion checking. Through this approach, JDart can either identify assertion violations, exhaustively explore all program paths, or reach resource limits during analysis.
@@
One of the key advantages of JDart is its capability to handle complex software systems, as exemplified by its development to analyze intricate NASA software. JDart’s architecture is modular, comprising two primary components: the executer and the explorer. The executer is responsible for executing the program and collecting symbolic constraints, leveraging the Java PathFinder framework [havelund2000model, visser2004test]. In contrast, the explorer determines the search and execution strategy employed during analysis.
@@
JDart was designed with two main objectives in mind. The first is to create a robust framework capable of managing industrial software challenges, such as handling crashes and addressing the constraints of constraint solving. The second objective is to establish a modular and extensible platform, allowing for the interchangeability of components, the use of various constraint solvers, and the implementation of multiple search strategies or termination conditions [jDart16].
@@
SAGE [Godefroid_Fuzz], developed by Microsoft, is an automated white-box fuzz testing tool widely utilized for detecting bugs and vulnerabilities in applications, such as those running on Windows. SAGE implements a technique known as Generational Search, which minimizes redundancy while maximizing the generation of new test cases. This systematic approach allows for effective analysis of the state space, enabling SAGE to manage large applications with substantial input sizes. Additionally, SAGE employs heuristics to enhance code coverage, further improving its effectiveness in identifying potential issues.
@@
Another tool developed by Godefroid is DART [Godefroid_DART], which focuses on automatic software testing using DSE. Building on DART’s capabilities, SMART [Godefroid_SMART] introduces the concept of compositional analysis, wherein composite functions are decomposed into atomic functions for individual analysis. The results of these analyses are then synthesized through the use of pre- and post-conditions.
@@
A key aspect that differentiates our research from previous work on DSE is our intended goal: to ascertain the semantic differences between executable component-and-connector architecture models. As noted in section˜1, semantic differencing operators have been developed for various modeling languages [MRR11a, MRR11b, DKMR19, DEKR19, Kau21], and component-and-connector structures, such as statechart systems (SCS) or automata, are no exception [BKRW17, BKRW19]. However, these approaches typically rely on translations to Büchi automata, which necessitate a finite state space and a defined input-output alphabet. Additionally, to compare entire architectures, syntactic composition of the automata is required, posing challenges when feedback loops are present. Our DSE-based approach, in contrast, circumvents these limitations.
@@ 4 Running Example
MontiArc [HRR12, Hab16, BKRW17a] is an architectural description language specifically designed for component-and-connector architectures. In this framework, components can either be atomic or consist of sub-architectures, allowing for a hierarchical organization of system elements. Communication between these components is achieved through a message-oriented approach, utilizing channels that connect output and input ports. The input-output behavior of atomic components is described by automata that are a variant of SCs [statechart, Rum16] based on the FOCUS formalism [BS12].
@@ 5.1 Major Design Decisions
The second reason for choosing DSE is that not all functions can be effectively analyzed using traditional symbolic execution. For instance, cryptographic functions, such as hash functions, cannot be evaluated through symbolic execution without undermining their security properties. In the case of DSE, these functions can be analyzed by utilizing concrete values for variables. By substituting concrete values, we can evaluate the function and consequently derive new input values for further analysis [Godefroid_SMART]. In summary, DSE was selected because it enables the analysis of functions that are not amenable to traditional symbolic execution and avoids exploring paths with unsatisfiable conditions.
@@
Z3 is an award-winning SMT solver developed by Microsoft. It is primarily used for predicate abstraction, advanced static checking, and test case generation. Interaction with Z3 can occur through the SMT-LIB format as well as various APIs [efficientZ3]. Our implementation utilizes the Java API for communication with the solver. Z3 supports multiple theories, including equality, arithmetic operations, and uninterpreted functions. Additionally, it generates a model containing concrete values for all defined constants within the formula. Z3 also includes numerous optimizations and tactics to enhance its performance [bjorner2019programming].
@@
Unlike other DSE engines that perform Dynamic Symbolic Execution by executing the program under test and collecting symbolic information without explicit code generation [Godefroid_Fuzz, jDart20, Sen], MontiArc employs a code generator that produces executable Java code [Hab16]. In this paper, we evaluate a DSE approach based on this existing code generator and extend it to generate symbolic code. We also discuss alternative methods, such as using an interpreter, in section˜7.
@@ 5.3 Computation of Semantic Differences Between MontiArc Models
Figure 7: Example of semantic difference between StudentVoteAlt and StudentVote, input length 3
When assessing the semantic difference between two models, it is crucial to consider certain specific scenarios. One such scenario arises when a component in the second model contains an automaton that is deemed partial concerning the possible inputs of the first model. In this case, the partial automaton will respond to any unknown input by ignoring it—meaning it neither produces a new output nor transitions to a different state [Rum96]. Consequently, the automaton retains its previous state and output. However, if the output ports are delayed, Null values may be generated. It is essential to address these cases within the automaton’s definition of the component behavior.
@@ 6.2 Results of the Evaluation
Figure 8: Results regarding the StudentVote based on input length 4
@@
In the subsequent analysis, we compute the semantic difference using the Path Coverage Controller with a garbage collection trigger. The results of the semantic difference analysis between StudentVote and StudentVoteAlt for varying input lengths are presented in table˜2.
@@
Table 2: Results of the computation of semantic differences between StudentVote and StudentVoteAlt
@@ 7 Discussion
Another avenue for enhancing the tool is through decomposition analysis of models, as introduced by Godefroid_SMART. This technique involves breaking down the program under analysis into individual functions. The results from analyzing these functions are then synthesized, taking into account their pre- and post-conditions, to evaluate composite programs. This methodology should be applicable to MontiArc models by examining their atomic components, which are already represented by atomic model artifacts. Implementing this approach will necessitate adjustments to the tool architecture.